OpenStudy (anonymous):

convrt 225 degrees to radians. explain how you found your answer.

OpenStudy (anonymous):

225/180=1.25 1.25=5/4 so 5pi/4 rads = 225degrees

OpenStudy (anonymous):

i tihnk thats how you do it

OpenStudy (lgbasallote):

to convert degrees to radians multiply your degree by \(\large \frac{\pi}{180}\)